The integrated luminosity for last week was 6.8 pb-1
Monday August 11
At 9:44 AM a quench at B1 ended shot setup.
The quench investigation eventually lead to the C49 separator, they eventually removed the reversing switch from the power supply.
The Tev access was completed by 6:30 PM, at which time the Tev turn-on ensued.
During the Tev startup a quench at BB occurred at 8:00 PM.
Quench recovery was completed by 9:30 PM, Tev turn-on then resumed.
Tuesday August 12
Stacking resumed and quench recovery began.
Tev was back on by 4:45 AM.
Another low beta quench occurred at 5:25 AM, experts began investigation.
The day shift began with the quench investigation continuing, stacking and beam to Miniboone.
At 9:45 AM Vacuum techs accessed the Tev to check vacuum in the quench related areas. No problem/cause were found.
Quench recovery was complete by 10:13 AM.
At 11:47 AM, D0 made a controlled access. The access was complete by 12:37 AM.
The Tev was back on by 1:00 PM.
Tev checkout and quench investigation continue while we wait for the stack to reach 120ma.
All studies were completed by 7:51 PM.
Shot setup for the Tev began at 8:08 PM, with store #2898 colliding by 9:46 PM with an initial average luminosity of 39.00e30.
We then returned to stacking.
Wednesday August 13
No problems the rest of the midnight shift.
Comments and Suggestions
|last modified 8/13/2003 email Fermilab|